OpenStudy (anonymous):

find vertex, line of symmetry and max/min -2x^2+2x+2 i got (1/2,5/2) x=1/2 min is 5/2

OpenStudy (anonymous):

yes you got it!

OpenStudy (anonymous):

axis is x=-b/2a=-2/(-4)=1/2 y-coordinate of vertex is y=f(1/2)=-1/2 + 1 +2 = 5/2 Vertex (1/2, 5/2) the vertex is a MAX though since a<0

OpenStudy (anonymous):

ahh ok, thanks!

OpenStudy (anonymous):

how do i get the perimeter of a rectangle

OpenStudy (anonymous):

sat is on you post now

OpenStudy (anonymous):

with a width of 8x^2cm and 12x^2y

OpenStudy (anonymous):

2L + 2W= P

OpenStudy (anonymous):

waht du u mean 2L+2W

OpenStudy (anonymous):